# The lower $p$-series of analytic pro-$p$ groups and Hausdorff dimension

### v2

@@ Abstract.
$\sup_{i\in\mathbb{N}}\;\Bigl{\lvert}\log_{p}\lvert G:P_{i}(G)\rvert-\Big{(}\sum_{k=1}^{d}\xi_{k}\Big{)}(i-1)\Bigr{\rvert}<\infty$
and such that the Hausdorff dimensions $\operatorname{hdim}_{G}^{\mathcal{L}}(H)$ of closed subgroups $H\leq_{\mathrm{c}}G$ with respect to $\mathcal{L}$ are confined to the set
@@ 1. Introduction
where we adopt the notation from [5, Sec. 1.2]. It is the fastest descending central series of $G$ with elementary abelian sections and it consists of open normal subgroups forming a base of neighbourhoods for the element $1$. The lower $p$-series plays a significant role in the study of pro-$p$ groups. For instance, $\bigoplus_{i=1}^{\infty}P_{i}(G)/P_{i+1}(G)$ naturally carries the structure of an $\mathbb{F}_{p}$-Lie algebra and can frequently be used to analyse the structure of $G$ or its finite quotients; compare with [13, Chap. VIII], [19, Chap. II] and [18, Sec. 5].
@@
In general, a filtration series of $L$, regarded as a $\mathbb{Z}_{p}$-lattice, is a descending chain $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ of open $\mathbb{Z}_{p}$-sublattices $L_{i}\leq_{\mathrm{o}}L$ such that $\bigcap_{i\in\mathbb{N}_{0}}L_{i}=\{0\}$. For $c\in\mathbb{N}_{0}$, we say that two filtration series $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ and $\mathcal{S}^{*}\colon L=L_{0}^{*}\geq L_{1}^{*}\geq\ldots$ are $c$-equivalent if for all $i\in\mathbb{N}$,
@@
Our aim is to show that the lower $p$-series of $L$ is equivalent to a filtration series that is descending very regularly, in a sense that we make precise now. Let $d=\dim_{\mathbb{Z}_{p}}L$ denote the dimension of the $\mathbb{Z}_{p}$-lattice $L$, and let $x_{1},\ldots,x_{d}$ be a $\mathbb{Z}_{p}$-basis for $L$. We say that a filtration series $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ of $L$ has a (vertical) stratification given by the frame $(x_{1},\ldots,x_{d})$ if there are positive real parameters $\xi_{1},\ldots,\xi_{d}$ with $\xi_{1}\leq\ldots\leq\xi_{d}$ such that $\mathcal{S}$ is equivalent to the filtration series
@@
The above result yields an analogous description of the lower $p$-series of a $p$-adic analytic pro-$p$ group. We recall that, for every $p$-adic analytic pro-$p$ group $G$, there exists $j\in\mathbb{N}$ such that for every $i\in\mathbb{N}$ with $i\geq j$ the group $P_{i}(G)$ is uniformly powerful; see [5, Prop. 3.9 and Thm. 4.2]. Moreover, every minimal generating system $x_{1},\ldots,x_{d}$ of a uniformly powerful pro-$p$ group $U$ yields a product decomposition of $U=\overline{\langle x_{1}\rangle}\cdots\overline{\langle x_{d}\rangle}$ into $d$ procyclic subgroups, which in turn gives a system of $p$-adic coordinates $U\to\mathbb{Z}_{p}^{\,d}$, $x_{1}^{\,a_{1}}\cdots x_{d}^{\,a_{d}}\mapsto(a_{1},\ldots,a_{d})$ of ‘the second kind’; see [5, Thm. 4.9]. The dimension of $G$, denoted by $\dim(G)$, coincides with the rank of any uniformly powerful open subgroup of $G$.
@@
One of our motivations for pinning down the lower $p$-series of $p$-adic analytic pro-$p$ groups was to settle some basic, but annoyingly outstanding questions about the Hausdorff spectra of such groups. The study and application of Hausdorff dimension in profinite groups is an ongoing affair; for instance, see [17] for an overview and [16, 7, 9, 10, 11, 12] for some recent developments. Let $G$ be an infinite finitely generated pro-$p$ group and let $\mathcal{S}\colon G_{i},\,i\in\mathbb{N}_{0}$, be a filtration series of $G$, that is, a descending chain $G=G_{0}\geq G_{1}\geq\ldots$ of open normal subgroups $G_{i}\trianglelefteq_{\mathrm{o}}G$ such that $\bigcap_{i\in\mathbb{N}_{0}}G_{i}=1$. In accordance with [1, 3], the Hausdorff dimension of a closed subgroup $H\leq_{\mathrm{c}}G$, with respect to $\mathcal{S}$, admits an algebraic interpretation as a logarithmic density:
@@
$\operatorname{hspec}^{\mathcal{S}}(G)=\{\operatorname{hdim}_{G}^{\mathcal{S}}(H)\mid H\leq_{\mathrm{c}}G\}\subseteq[0,1]$
@@
Even for well-behaved groups, such as $p$-adic analytic pro-$p$ groups $G$, not only the Hausdorff dimension of individual subgroups, but also qualitative features of the Hausdorff spectrum of $G$ are known to be sensitive to the choice of $\mathcal{S}$; see [17, Thm. 1.3]. In our setting, where $G$ is a finitely generated pro-$p$ group, the group-theoretically relevant choices of $\mathcal{S}$ are primarily: the $p$-power series $\mathcal{P}$, the iterated $p$-power series $\mathcal{P}^{*}$, the Frattini series $\mathcal{F}$, the dimension subgroup (or Zassenhaus) series $\mathcal{D}$ and, last but not least, the lower $p$-series $\mathcal{L}\colon P_{i}(G),\,i\in\mathbb{N}$. It was shown in [17, Prop. 1.5] that, if $G$ is a $p$-adic analytic pro-$p$ group, then the Hausdorff dimensions of $H\leq_{\mathrm{c}}G$ with respect to the series $\mathcal{P},\mathcal{F},\mathcal{D}$ coincide. Furthermore, [3, Thm. 1.1] yields, for $\mathcal{S}$ any one of these series $\mathcal{P},\mathcal{F},\mathcal{D}$, that the spectrum
$\operatorname{hspec}^{\mathcal{S}}(G)=\Big{\{}\nicefrac{{\displaystyle\dim(H)}}{{\displaystyle\dim(G)}}\mid H\leq_{\mathrm{c}}G\Big{\}}$
@@
It was seen in [17] that, contrary to indications in [3], the Hausdorff spectra of $p$-adic analytic pro-$p$ groups $G$ with respect to the lower $p$-series $\mathcal{L}$ follow a different pattern. Concretely, [17, Example 4.1] provides a family of $p$-adic analytic pro-$p$ groups $G$ such that $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert/\dim(G)$ is unbounded as $\dim(G)\to\infty$. It remained open whether $\operatorname{hspec}^{\mathcal{L}}(G)$ could even be infinite for some $p$-adic analytic groups $G$. Theorem 1.1 allows us to establish that $\operatorname{hspec}^{\mathcal{L}}(G)$ is in fact finite for every $p$-adic analytic pro-$p$ group $G$ and to put in place further restrictions.
@@
Theorem 1.5 triggers more refined questions about Hausdorff spectra with respect to the lower $p$-series. We address some of these in a separate paper. For instance, one outstanding problem is to understand what renders the Hausdorff spectrum of a finitely generated pro-$p$ group finite. Using Lie-theoretic tools we show that non-abelian finite direct products of free pro-$p$ groups have full Hausdorff spectrum $[0,1]$ with respect to the lower $p$-series, in analogy to [7].
@@ 2. Stratifications of the lower $p$-series
Let $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ be a filtration series of a $\mathbb{Z}_{p}$-lattice $L$ which admits a stratification. Then any two stratifications of $\mathcal{S}$ have the same growth rate.
@@
The following auxiliary result is a variation of [14, Prop. 4.3] and can be proved in the same way, using compactness.
@@
Denote by $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, the terms of the lower $p$-series of $L$, and for $i\in\mathbb{N}_{0}$ put $u_{i}=u_{L}(L_{i})$. By Proposition 2.2 the $\mathbb{Z}_{p}G$-module $L$ has finite rigidity $r=r(L)$. Consequently, the lower $p$-series of $L$ is $r$-equivalent to the filtration series $\mathcal{S}\colon L\geq p^{u_{1}}L\geq p^{u_{2}}L\geq\ldots$, and it suffices to show that $\mathcal{S}$ is equivalent to a series of the form $p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$, for suitable $\xi\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$.
[…]
@@
defined by the dictum $p^{u_{i}}x_{k}\mapsto z_{k}$ for $1\leq k\leq d$. For each $i\in\mathbb{N}$, we have $\ell_{L}(L_{i})\leq u_{i}+r$, hence $p^{u_{i}+r}L\leq L_{i}\leq p^{u_{i}}L$ by the definition of $u_{i}$ and $r$, and we consider $R_{i}=L_{i}\varphi_{i}$.
@@
For a contradiction, assume that $M$ is not $G$-invariant, i.e. not a $\mathbb{Z}_{p}G$-submodule. Then there are elements $\alpha_{1},\ldots,\alpha_{e}\in\mathbb{Z}_{p}G$ and $c_{2}\in\mathbb{N}$ such that $y=\sum_{k=1}^{e}x_{k}\alpha_{k}\in N\smallsetminus p^{c_{2}}N$. For every $i\in\mathbb{N}$, we deduce from $\xi_{e}=\max\{\xi_{1},\ldots,\xi_{e}\}$ that $p^{\lfloor i\xi_{e}\rfloor+c_{1}}M\subseteq L_{i}$ and hence
@@
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}y_{k}\leq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$
@@
We are now in a position to prove Theorem 1.1.
@@
Let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a $G$-action, for a finitely generated pro-$p$ group $G$. In view of Lemma 2.1, Proposition 2.4 and Corollary 2.5, it remains to prove that the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ admits a stratification with rational growth rates in $[\nicefrac{{1}}{{d}},1]$, where $d=\dim_{\mathbb{Z}_{p}}L$.
We argue by induction on $d$. Of course, the claim holds trivially if $L=\{0\}$, and we now suppose that $d\geq 1$. Let $N$ be a $\mathbb{Z}_{p}G$-submodule of $L$ such that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is a simple $\mathbb{Q}_{p}G$-module. Choose $n\in\mathbb{N}$ such that $p^{n+1}L\cap N\subseteq pN$. Then $\tilde{L}=p^{n}L+N$ is a $G$-invariant open $\mathbb{Z}_{p}$-sublattice of $L$, and $N$ is a direct summand of the $\mathbb{Z}_{p}$-lattice $\tilde{L}$. By Lemma 2.6, there is no harm in replacing $L$ with $\tilde{L}$ and we may thus suppose that $L$ itself decomposes as $L=M\oplus N$ for a $\mathbb{Z}_{p}$-sublattice $M$.
@@
Now suppose that $N$ has no almost complement in $L$. Let us fix some additional notation. We write
@@
for the lower $p$-series. Put $e=\dim_{\mathbb{Z}_{p}}M$. By Proposition 2.3 there is $\eta\in[\nicefrac{{1}}{{(d-e)}},1]\cap\mathbb{Q}$ such that the lower $p$-series of $N$ is equivalent to the series $p^{\lfloor i\eta\rfloor}N$, $i\in\mathbb{N}_{0}$. By induction on the dimension, the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L/N$, which is $(L_{i}+N)/N$, $i\in\mathbb{N}$, admits a stratification, based on a frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$, with growth rates $\xi_{1}\leq\ldots\leq\xi_{e}$ in $[\nicefrac{{1}}{{e}},1]\cap\mathbb{Q}$. Set $\xi_{0}=0$ for notational convenience.
We choose representatives $x_{1},\ldots,x_{e}\in M$ for the cosets $\mathbf{x}_{1},\ldots,\mathbf{x}_{e}\in L/N$, and pick $\tilde{e}\in\{0,1,\ldots,e-1\}$ as large as possible subject to the conditions:
@@
In fact, we arrange that among all possible choices – for $M$ among all possible almost complements for $N$ in $L$ (with a subsequent reduction to $L=M\oplus N$), for the stratification based on the frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$ and for the representatives $x_{1},\ldots,x_{e}$ in $L$ – we make one that renders $\tilde{e}$ as large as possible. In addition we choose $\hat{e}\in\{\tilde{e}+1,\ldots,e\}$ maximal with $\xi_{\hat{e}}=\xi_{\tilde{e}+1}$, and for short we put
$\vartheta=\xi_{\tilde{e}+1}=\xi_{\tilde{e}+2}=\ldots=\xi_{\hat{e}}\qquad\text{and}\qquad T=\bigoplus_{k=\tilde{e}+1}^{\hat{e}}\mathbb{Z}_{p}x_{k}.$

### v3

@@ Abstract.
$\sup_{i\in\mathbb{N}}\;\Bigl{\lvert}\log_{p}\lvert G:P_{i}(G)\rvert-\Big{(}\sum_{k=1}^{d}\xi_{k}\Big{)}(i-1)\Bigr{\rvert}$$\infty$
and such that the Hausdorff dimensions $\operatorname{hdim}_{G}^{\mathcal{L}}(H)$ of closed subgroups $H\subseteq_{\mathrm{c}}G$ with respect to $\mathcal{L}$ are confined to the set
@@ 1. Introduction
where we adopt the notation from [5, Sec. 1.2]. It is the fastest descending central series of $G$ with elementary abelian sections and it consists of open normal subgroups forming a base of neighbourhoods for the element $1$. The lower $p$-series plays a significant role in the study of pro-$p$ groups. For instance, $\bigoplus_{i=1}^{\infty}P_{i}(G)/P_{i+1}(G)$ naturally carries the structure of an $\mathbb{F}_{p}$-Lie algebra and can frequently be used to analyse the structure of $G$ or its finite quotients; compare with [14, Chap. VIII], [20, Chap. II] and [19, Sec. 5].
@@
In general, a filtration series of $L$, regarded as a $\mathbb{Z}_{p}$-lattice, is a descending chain $\mathcal{S}\colon L=L_{0}\supseteq L_{1}\supseteq\ldots$ of open $\mathbb{Z}_{p}$-sublattices $L_{i}\subseteq_{\mathrm{o}}L$ such that $\bigcap_{i\in\mathbb{N}_{0}}L_{i}=\{0\}$. For $c\in\mathbb{N}_{0}$, we say that two filtration series $\mathcal{S}\colon L=L_{0}\supseteq L_{1}\supseteq\ldots$ and $\mathcal{S}^{*}\colon L=L_{0}^{*}\supseteq L_{1}^{*}\supseteq\ldots$ are $c$-equivalent if for all $i\in\mathbb{N}$,
@@
Our aim is to show that the lower $p$-series of $L$ is equivalent to a filtration series that is descending very regularly, in a sense that we make precise now. Let $d=\dim_{\mathbb{Z}_{p}}L$ denote the dimension of the $\mathbb{Z}_{p}$-lattice $L$, and let $x_{1},\ldots,x_{d}$ be a $\mathbb{Z}_{p}$-basis for $L$. We say that a filtration series $\mathcal{S}\colon L=L_{0}\supseteq L_{1}\supseteq\ldots$ of $L$ has a (vertical) stratification given by the frame $(x_{1},\ldots,x_{d})$ if there are positive real parameters $\xi_{1},\ldots,\xi_{d}$ with $\xi_{1}\leq\ldots\leq\xi_{d}$ such that $\mathcal{S}$ is equivalent to the filtration series
@@
The proof of Theorem 1.1 is relatively straightforward in the special case where $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a completely reducible $\mathbb{Q}_{p}G$-module; see Remark 2.7. But the general setting requires additional care. In essence, our analysis gives rise to an inductive procedure for computing the growth rate parameters; see Remark 2.8.
By means of Lie-theoretic considerations, Theorem 1.1 yields an analogous description of the lower $p$-series of a $p$-adic analytic pro-$p$ group. We recall that, for every $p$-adic analytic pro-$p$ group $G$, there exists $j\in\mathbb{N}$ such that for every $i\in\mathbb{N}$ with $i\geq j$ the group $P_{i}(G)$ is uniformly powerful; see [5, Prop. 3.9 and Thm. 4.2]. Moreover, every minimal generating system $x_{1},\ldots,x_{d}$ of a uniformly powerful pro-$p$ group $U$ yields a product decomposition of $U=\overline{\langle x_{1}\rangle}\cdots\overline{\langle x_{d}\rangle}$ into $d$ procyclic subgroups, which in turn gives a system of $p$-adic coordinates $U\to\mathbb{Z}_{p}^{\,d}$, $x_{1}^{\,a_{1}}\cdots x_{d}^{\,a_{d}}\mapsto(a_{1},\ldots,a_{d})$ of ‘the second kind’; see [5, Thm. 4.9]. The dimension of $G$, denoted by $\dim(G)$, coincides with the rank of any uniformly powerful open subgroup of $G$.
@@
One of our motivations for pinning down the lower $p$-series of $p$-adic analytic pro-$p$ groups was to settle some basic, but annoyingly outstanding questions about the Hausdorff spectra of such groups. The study and application of Hausdorff dimension in profinite groups is an ongoing affair; for instance, see [18] for an overview and [17, 7, 9, 10, 12, 13] for some recent developments. Let $G$ be an infinite finitely generated pro-$p$ group and let $\mathcal{S}\colon G_{i},\,i\in\mathbb{N}_{0}$, be a filtration series of $G$, that is, a descending chain $G=G_{0}\supseteq G_{1}\supseteq\ldots$ of open normal subgroups $G_{i}\trianglelefteq_{\mathrm{o}}G$ such that $\bigcap_{i\in\mathbb{N}_{0}}G_{i}=1$. In accordance with [1, 3], the Hausdorff dimension of a closed subgroup $H\subseteq_{\mathrm{c}}G$, with respect to $\mathcal{S}$, admits an algebraic interpretation as a logarithmic density:
@@
$\operatorname{hspec}^{\mathcal{S}}(G)=\{\operatorname{hdim}_{G}^{\mathcal{S}}(H)\mid H\subseteq_{\mathrm{c}}G\text{ a subgroup}\}\subseteq[0,1]$
@@
Even for well-behaved groups, such as $p$-adic analytic pro-$p$ groups $G$, not only the Hausdorff dimension of individual subgroups, but also qualitative features of the Hausdorff spectrum of $G$ are known to be sensitive to the choice of $\mathcal{S}$; see [18, Thm. 1.3]. In our setting, where $G$ is a finitely generated pro-$p$ group, the group-theoretically relevant choices of $\mathcal{S}$ are primarily: the $p$-power series $\mathcal{P}$, the iterated $p$-power series $\mathcal{P}^{*}$, the Frattini series $\mathcal{F}$, the dimension subgroup (or Zassenhaus) series $\mathcal{D}$ and, last but not least, the lower $p$-series $\mathcal{L}\colon P_{i}(G),\,i\in\mathbb{N}$. It was shown in [18, Prop. 1.5] that, if $G$ is a $p$-adic analytic pro-$p$ group, then the Hausdorff dimensions of a subgroup $H\subseteq_{\mathrm{c}}G$ with respect to the series $\mathcal{P},\mathcal{F},\mathcal{D}$ coincide. Furthermore, [3, Thm. 1.1] yields, for $\mathcal{S}$ any one of these series $\mathcal{P},\mathcal{F},\mathcal{D}$, that the spectrum
$\operatorname{hspec}^{\mathcal{S}}(G)=\Big{\{}\nicefrac{{\displaystyle\dim(H)}}{{\displaystyle\dim(G)}}\mid H\subseteq_{\mathrm{c}}G\text{ a subgroup}\Big{\}}$
@@
It was seen in [18] that, contrary to indications in [3], the Hausdorff spectra of $p$-adic analytic pro-$p$ groups $G$ with respect to the lower $p$-series $\mathcal{L}$ follow a different pattern. Concretely, [18, Example 4.1] provides a family of $p$-adic analytic pro-$p$ groups $G$ such that $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert/\dim(G)$ is unbounded as $\dim(G)\to\infty$. It remained open whether $\operatorname{hspec}^{\mathcal{L}}(G)$ could even be infinite for some $p$-adic analytic groups $G$. Theorem 1.1 allows us to establish that $\operatorname{hspec}^{\mathcal{L}}(G)$ is in fact finite for every $p$-adic analytic pro-$p$ group $G$ and to put in place further restrictions.
@@
Theorem 1.5 triggers more refined questions about Hausdorff spectra with respect to the lower $p$-series. We address some of these in a separate paper [11]. For instance, one outstanding problem is to understand what renders the Hausdorff spectrum of a finitely generated pro-$p$ group finite. Using Lie-theoretic tools we show that non-abelian finite direct products of free pro-$p$ groups have full Hausdorff spectrum $[0,1]$ with respect to the lower $p$-series, in analogy to [7].
@@
Acknowledgements. We thank both referees for very useful comments and suggestions.
@@ 2. Stratifications of the lower $p$-series
Let $\mathcal{S}\colon L=L_{0}\supseteq L_{1}\supseteq\ldots$ be a filtration series of a $\mathbb{Z}_{p}$-lattice $L$ which admits a stratification. Then any two stratifications of $\mathcal{S}$ have the same growth rate.
@@
The following auxiliary result is a variation of [15, Prop. 4.3] and can be proved in the same way, using compactness.
@@
Denote by $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, the terms of the lower $p$-series of $L$, and for $i\in\mathbb{N}_{0}$ put $u_{i}=u_{L}(L_{i})$. By Proposition 2.2 the $\mathbb{Z}_{p}G$-module $L$ has finite rigidity $r=r(L)$. Consequently, the lower $p$-series of $L$ is $r$-equivalent to the filtration series $\mathcal{S}\colon L\supseteq p^{u_{1}}L\supseteq p^{u_{2}}L\supseteq\ldots$, and it suffices to show that $\mathcal{S}$ is equivalent to a series of the form $p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$, for suitable $\xi\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$.
Write $L=\bigoplus_{k=1}^{d}\mathbb{Z}_{p}x_{k}$ and choose, as a reference group, a homogeneous abelian group of exponent $p^{r}$ on $d$ generators, $R=\langle z_{1},\ldots,z_{d}\rangle\cong C_{p^{r}}\times\overset{d}{\ldots}\times C_{p^{r}}$. For each $i\in\mathbb{N}$ consider the epimorphism
@@
defined by the dictum $p^{u_{i}}x_{k}\mapsto z_{k}$ for $1\leq k\leq d$. For each $i\in\mathbb{N}$, we have $\ell_{L}(L_{i})\leq u_{i}+r$, hence $p^{u_{i}+r}L\subseteq L_{i}\subseteq p^{u_{i}}L$ by the definition of $u_{i}$ and $r$, and we consider $R_{i}=L_{i}\varphi_{i}$.
@@
For a contradiction, assume that $M$ is not $G$-invariant, i.e. not a $\mathbb{Z}_{p}G$-submodule. Then there are $\alpha_{1},\ldots,\alpha_{e}\in\mathbb{Z}_{p}G$ such that $y=\sum_{k=1}^{e}x_{k}\alpha_{k}\in N\smallsetminus\{1\}$. Choose $c_{2}\in\mathbb{N}$ such that $y\in N\smallsetminus p^{c_{2}}N$. For every $i\in\mathbb{N}$, we deduce from $\xi_{e}=\max\{\xi_{1},\ldots,\xi_{e}\}$ that $p^{\lfloor i\xi_{e}\rfloor+c_{1}}M\subseteq L_{i}$ and hence
@@
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}y_{k}\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$
@@
We are now in a position to prove Theorem 1.1, but it is worth to comment first on a special case.
Remark 2.7.
Proposition 2.3 and Lemma 2.6 already yield a direct proof of Theorem 1.1 in the special case where $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a completely reducible $\mathbb{Q}_{p}G$-module. In this context we recall the following fact: if the pro-$p$ group $G$ is an open compact subgroup of a semisimple $p$-adic group, i.e. the group $\mathbf{G}(\mathbb{Q}_{p})$ of $\mathbb{Q}_{p}$-rational points of a semisimple linear algebraic group $\mathbf{G}$ over $\mathbb{Q}_{p}$, then every $\mathbb{Q}_{p}G$-module $V$ with $\dim_{\mathbb{Q}_{p}}(V)<\infty$ is completely reducible. Indeed, for every open subgroup $H$ of $G$, the $\mathbb{Q}_{p}G$-module $V$ is completely reducible if and only if it is completely reducible as a $\mathbb{Q}_{p}H$-module, essentially by the standard Clifford theorem and Maschke theorem. Hence, by passing to an open subgroup of $G$, we may assume that $G$ is uniformly powerful and form the $\mathbb{Q}_{p}$-Lie algebra $\mathcal{L}_{G}=\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}\log(G)$ associated to $G$. In this set-up, it is easily seen that $\mathbb{Q}_{p}G$-submodules of $V$ are the same as $\mathcal{L}_{G}$-submodules of $V$. Since the Lie algebra $\mathcal{L}_{G}$ is semisimple of characteristic zero, Weyl’s theorem shows that $V$ is completely reducible.
@@
Let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a $G$-action, for a finitely generated pro-$p$ group $G$. In view of Lemma 2.1, Proposition 2.4 and Corollary 2.5, it remains to prove that the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ admits a stratification with rational growth rates in $\nicefrac{{1}}{{d}},1]$, where $d=\dim_{\mathbb{Z}_{p}}L$.
We argue by induction on $d$. Of course, the claim holds trivially if $L=\{0\}$, and we now suppose that $d\geq 1$. By intersecting a simple $\mathbb{Q}_{p}G$-submodule of $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ with the original lattice $L$, we obtain a $\mathbb{Z}_{p}G$-submodule $N$ of $L$ such that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is a simple $\mathbb{Q}_{p}G$-module. Choose $n\in\mathbb{N}$ such that $p^{n+1}L\cap N\subseteq pN$. Then $\tilde{L}=p^{n}L+N$ is a $G$-invariant open $\mathbb{Z}_{p}$-sublattice of $L$, and $N$ is a direct summand of the $\mathbb{Z}_{p}$-lattice $\tilde{L}$. By Lemma 2.6, there is no harm in replacing $L$ with $\tilde{L}$ and we may thus suppose that $L$ itself decomposes as $L=M\oplus N$ for a $\mathbb{Z}_{p}$-sublattice $M$.
@@
Now suppose that $N$ has no $G$-invariant almost complement in $L$. Let us fix some additional notation. We write
@@
for the lower $p$-series. Put $e=\dim_{\mathbb{Z}_{p}}M$. By Proposition 2.3 there is $\eta\in[\nicefrac{{1}}{{(d-e)}},1]\cap\mathbb{Q}$ such that the lower $p$-series of $N$ is equivalent to the series $p^{\lfloor i\eta\rfloor}N$, $i\in\mathbb{N}_{0}$. By induction on the dimension, the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L/N$, which is $(L_{i}+N)/N$, $i\in\mathbb{N}$, admits a stratification, based on a frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$, with growth rates $\xi_{1}\leq\ldots\leq\xi_{e}$ in $\nicefrac{{1}}{{e}},1]\cap\mathbb{Q}$. Set $\xi_{0}=0$ for notational convenience.
We choose representatives $x_{1},\ldots,x_{e}\in M$ for the cosets $\mathbf{x}_{1},\ldots,\mathbf{x}_{e}\in L/N$, and pick $\widetilde{e}\in\{0,1,\ldots,e-1\}$ as large as possible subject to the conditions:
$\xi_{\widetilde{e}}<\xi_{\widetilde{e}+1}\qquad\text{and}\qquad\tilde{M}=\bigoplus_{k=1}^{\widetilde{e}}\mathbb{Z}_{p}x_{k}\quad\text{is $G$-invariant.}$
@@
$\vartheta=\xi_{\widetilde{e}+1}=\xi_{\widetilde{e}+2}=\ldots=\xi_{\widehat{e}}\qquad\text{and}\qquad T=\bigoplus_{k=\widetilde{e}+1}^{\widehat{e}}\mathbb{Z}_{p}x_{k}.$